40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
Message ID: 16875     Entry time: Wed May 25 17:34:47 2022     In reply to: 16874     Reply to this: 16877
Author: yuta 
Type: Configuration 
Category: BHD 
Subject: IFO recovery - IFO alignment 

IFO aligned to maximize flashings, except for GRY and LO-AS.

What we did:
 0. After recovering IMC, C1:IOO-MC_TRANS_SUM was ~1300 with C1:IOO-MC_RFPD_DCMON of ~0.11 (~10% better than what we had during vent). Xarm and Yarm was already flashing and could see the beam at AS and POP cameras.
 1. Aligned ETMX and ITMX to green X input beam to maximize C1:ALS-TRX_OUT, to ~0.19.
 2. Aligned TT2-PR3 to get C1:SUS-ETMX_TRX_OUT flashing at 0.09 at max
 3. Aligned ITMY to have nice POP blinking of MICH at POP camera
 4. Aligned ETMY-PR3 to have C1:SUS-ETMX_TRX_OUT flashing at 0.06 at max
 5. Misaligned ITMY (with +2 in C1:SUS-ITMY_PIT_COMM), and aligned PRM to have PRX (PRM-ITMX cavity) flashing at C1:LSC-ASDC_IN1 at ~20 (offset -70) at max
 6. Misaligned PRM, and aligned SRM to have SRX (SRM-ITMX cavity) flashing at C1:LSC-ASDC_IN1 at ~20 (offset -70) at max
 7. Restored all the alignment. ITMY didn't quite come back, so I need to tweak the alignement to maximize TRY flashing.

Result:
Current alignment is as attached. IR beam at AS, REFL, MCR and green beam at GTRX cameras all seem slightly to the left from monitors, but looks as it was before the pump down.yes GTRY is still clipped, but green Y locks stably. Oplevs were not so useful to recover the alignment. ETMX/Y oplevs did not drifted too much probably because we don't have in-vac steering mirrors.

Next:
 - Tweak alignment of green Y input to follow Yarm
 - Do LO-AS alignment
 - REFL DC is not receiving beam. Re-alignment necessary
 - Oplev centering
 - BHD PDs need to be replaced to lower gain PDs and need to be connected to CDS

Attachment 1: Screenshot_2022-05-25_17-47-57.png  385 kB  Uploaded Wed May 25 18:48:11 2022  | Hide | Hide all
Screenshot_2022-05-25_17-47-57.png
ELOG V3.1.3-